# Multimodal Laryngoscopic Video Analysis for Assisted Diagnosis of Vocal Fold Paralysis

### v3

@@ 1 Introduction
Vocal fold Paralysis (VFP) is a condition where one of the vocal folds fails to move properly, leading to voice changes, difficulty swallowing, and potential breathing problems [1, 2]. VFP can result from nerve damage due to surgery, injury, infection, or tumors, significantly impacting a patient’s quality of life [3]. Accurate diagnosis of VFP is crucial, as it informs the appropriate medical or surgical intervention, which can restore vocal function, improve airway protection, and enhance overall patient outcomes. Clinicians often use laryngeal videostroboscopy to check the vocal fold vibration in details. Laryngeal videostroboscopy is a specialized diagnostic tool used to evaluate the function of the vocal folds and the larynx. The stroboscopic component is important because it allows for visualization of vocal fold vibrations, which are often too rapid to be observed directly by doctors [4]. Stroboscopy uses a light source that flashes at a slightly different frequency than the vocal fold vibration. This creates a slow-motion effect, enabling the clinician to see the individual phases of vocal fold vibration in great details.
With the advent of the artificial intelligence, deep learning methods are developed to extract useful parameters [5] to assist clinicians track the motion of vocal folds [6, 7, 8, 9, 10], and even make predictions [11, 12] or classifications [13] to help clinicians perform diagnosis. However, the diagnosis of vocal folds requires to inspect the complete phonation cycles of patients. Previous works default to using video frames or images that already contain the vocal folds, but in most of the time, the raw recordings from the endoscopic inspection often carry useless information other than patient’s phonation cycles. For instance, no phonation cycles occur in the beginning of the inspection, as the laryngoscope is still finding the vocal folds. As a result, manual labeling and selection of video frames from raw endoscopic videos are required, making the diagnosis process time-consuming and inconvenient for clinicians.
@@
In this article, we introduce the Multimodal Laryngoscopic Video Analyzing System (MLVAS), a novel framework that leverages both audio and visual modalities to automatically process and analyze raw laryngeal videostroboscopic recordings. The proposed system comprises a multimodal front-end module, a feature extraction module, and a multimodal back-end classification module for detecting VFP. The front-end module identifies and extracts video segments that captures complete phonation cycles with clear vocal folds’ movement from raw endoscopic laryngeal videos. Subsequently, MLVAS extracts both audio and video features from these segments. By leveraging these complementary modalities, the system employs a classifier in the back-end for VFP detection. The primary contributions of this work are as follows:
@@
    The system introduces Left and Right Vocal Fold Dynamics (LVFDyn and RVFDyn) to reflect the vocal fold’s activity, enabling precise diagnosis of left- or right-sided VFP. Experimental findings highlight its effectiveness in diagnosing Unilateral Vocal Fold Paralysis (UVFP).
@@ 2.1 System Design
2.2 Audio Processing Module
To get a better examination of the status of the vocal folds, patients are often pronouncing ”e” (although asked to pronounce ”a”) while the clinicians examine the larynx using the endoscope. In this way, the vibration of the vocal folds can be easily detected by the clinicians. Hence, by capturing the audio segments that pronounce ”e”, phonation cycles can be detected. To accomplish this, we developed a KWS model that is used to detect audio keywords in a sentence, such as ”Hey Siri”, ”Ok Google”, and in our case, the word ”e”.
The overall pipeline is illustrated in Figure 2. Initially, the input audio is transformed into a STFT spectrogram, converting the time-domain signal into a time-frequency representation for more effective analysis. The spectrogram is then segmented into chunks along the time axis, with each chunk containing a fixed number of frames. During the training phase, these chunks are randomly selected from each audio clip to ensure a diverse set of training samples. This method enhances the model’s generalization capabilities by incorporating spectrograms from various audio clips into a robust batch of training data.
@@ 2.2 Audio Processing Module
During inference, the input spectrogram is sliced into chunks using a sliding window mechanism. Each chunk is processed by the trained KWS model to generate decision results, as depicted in Figure 2. This KWS model can reliably detects the ”e” vocalization, providing critical prior knowledge for the subsequent analysis.
@@ 2.3 Video Processing Module
Although the audio KWS model is employed to detect the vocalization video segment of the phonation ”e”, these detected video segments may not always successfully capture the vocal folds and glottis on the visual side. To address this issue, we train a vocal fold detection model, utilizing the famous object detection model structure YOLO-v5 [38] on the BAGLS dataset [39], to further refine the time masks.
As depicted in Figure 1, our pipeline integrates a strobing video extraction module alongside the vocal fold image object detection module. This module is utilized to select the strobing segments of laryngeal videos by analyzing hue, saturation, and value (HSV) of video frames. The HSV parameters represent type, intensity, and brightness of colors, respectively. Transitions marked by empty frames typically occur at the beginning and end of strobing segments. We employ a unit step function to identify and mark all empty frames with a zero value in color (illustrated by the yellow line in Figure 3). This technique segments the video into several continuous non-empty frame sequences. Within these non-empty segments, we examine the variations in HSV values. The variations is calculated as the total number of changes within a time range. Due to the rapid color-changing characteristic of strobing videos (as shown in Figure 3), we are able to identify the strobing segments by calculating the frequency of HSV fluctuations within each non-empty segment, selecting the segment with the highest number of fluctuations. Suppose $v_{t-1},v_{t},v_{t+1}$ stands for the value at time step $t-1,t$ and $t+1$ respectively. Then, the frequency of HSV fluctuation within a time range of $T=[t_{0},t_{n}]$ is calculated as follows.
$F_{T}=\sum_{t=1}^{n-1}\frac{(v_{t}-v_{t-1})*(v_{t+1}-v_{t})}{|(v_{t}-v_{t-1})*(v_{t+1}-v_{t})|}$
@@ 4.2 Diffusion Model-based Refinement
Algorithm 1 Angle Deviation Extraction Algorithm
0:  Glottis mask vertices’ coordinates $U(x_{1},y_{1})$, $D(x_{2},y_{2})$, $L(x_{3},y_{3})$, $R(x_{4},y_{4})$
0:  Glottal angles $\{\angle L_{k}D^{\prime}C\}_{k\in[1,N-1]}$, $\{\angle CD^{\prime}R_{k}\}_{k\in[1,N-1]}$
1:  Compute the center $C(x_{c},y_{c})$ from $U,D,L,R$
2:  Compute the function $f(x)$ passing through $C,D$
3:  Find $N$ equidistant points $C_{1},C_{2},...,C_{N-1}$ of $f(x)$ on the line segment intercepted by the glottal mask
4:  for $k=1$ to $N-1$ do
5:     Compute the function $f^{\prime}_{k}(x)$ orthogonal to $f(x)$ passing through $C_{k}$
6:     Compute intersection points $L_{k},R_{k}$ between $f_{k}^{\prime}(x)$ and the glottis mask
7:  end for
8:  Rotate the coordinate system by $\gamma$ degree to make $f(x)$ vertically aligned
9:  Approximate a quadratic curve $q_{\gamma}(x)$ using rotated intersection points $\{L_{\gamma,k}\}_{k\in[1,N-1]}$ and $\{R_{\gamma,k}\}_{k\in[1,N-1]}$ in the rotated coordinate system
10:  Find the vertex point $D_{q}$ of $q_{\gamma}(x)$ and map $D_{q}$ to the original coordinate system
11:  Connect $C$ and $D_{q}$ to form the corrected middle line $f^{*}(x)$ and intersect the glottis mask at $D^{\prime}$
12:  Connect $\{L_{k}\}_{k\in[1,N-1]}$ and $\{R_{k}\}_{k\in[1,N-1]}$ with $D^{\prime}$ to obtain glottal angles
Figure 5: The workflow of computing the angle deviation of left and right vocal folds. Step (a) get the center point $C$ and bottom point $D$. Step (b), (c) and (d) fit the outline of vocal folds with a quadratic function, refining the bottom point to $D^{\prime}$. As a result, step (e) shows the refined midline $CD^{\prime}$. Finally, the angle deviation for both vocal folds can be computed.
@@
4.3 Vocal Fold Dynamics (VFDyn) Extraction with Quadratic Fitting
To diagnose the laryngeal paralysis in a more detailed way, instead of AGA or GAW that measures the both left and right vocal folds, we evaluate the angle deviation for each vocal fold. As a result, we can provide a more comprehensive diagnosis of UVFP. The procedure for extracting glottal angles from a segmented glottis involves a systematic series of steps, detailed in Algorithm 1, with visual representations of the intermediate results provided in Figure 5.
Initially, the algorithm acquires the coordinates of the top, bottom, left, and right vertices from the glottis mask, computing their center point denoted as $U(x_{1},y_{1})$, $D(x_{2},y_{2})$, $L(x_{3},y_{3})$, $R(x_{4},y_{4})$, and $C(x_{c},y_{c})$ respectively (see Figure 5(a)). Subsequently, a line connecting points $C$ and $D$ is established, and a linear function $f(x)$ passing through $C$ and $D$ is computed.
Equidistant points $C_{1},C_{2},...,C_{N-1}$ are then positioned along the line segment intercepted by the glottal mask. For each $C_{k}$, orthogonal functions $f^{\prime}_{k}(x)$ to $f(x)$ are calculated, with their intersection points $L_{k}$ and $R_{k}$ with the glottis mask determined for $k\in[1,N-1]$ (see Figure 5(b)).
To ensure uniformity, the coordinate system is rotated by an angle $\gamma$, aligning all intersection points along the y-axis after rotation, denoted as $L_{\gamma,k}$ and $R_{\gamma,k}$. Leveraging these points, a quadratic curve $q_{\gamma}(x)$ is approximated in the rotated coordinate system (see Figure 5(c). The lowest point $D_{q}$ of $q_{\gamma}(x)$ is identified and mapped back to the original coordinate system (see Figure 5(d)).
Subsequently, a calibrated middle line $f^{*}(x)$ connecting points $C$ and $D_{q}$ is established, intersecting the glottis mask at point $D^{\prime}$. Finally, glottal angles are derived by connecting points $\{L_{k}\}_{k\in[1,N-1]}$ and $\{R_{k}\}_{k\in[1,N-1]}$ with $D^{\prime}$, yielding $\{\angle L_{k}D^{\prime}C\}_{k\in[1,N-1]}$ and $\{\angle CD^{\prime}R_{k}\}_{k\in[1,N-1]}$ respectively.
@@ 6.3 Keyword Spotting Model
Since the KWS model outputs a posterior score rather than a direct decision value, a threshold is necessary to determine the final result. Consequently, the threshold is selected based on the highest F-score, a balanced metric that considers both precision and recall, providing a comprehensive measure of classification performance. In MLVAS, we pick a threshold of 0.38 as the classification boundary. If the posterior score output by the KWS model is higher than this value, the input chunk is regarded as ”e”. Otherwise, it is treated as other voices. Figure 7(a) and Figure 7(b) display the confusion matrix computed on the SYSU-B-test and SYSU-A dataset respectively. The accuracy, precision, recall and F-score are listed in Table 4. As is shown in Table 4, the performance of our audio KWS model are all above 90% on both the SYSU-A and SYSU-B.
@@ 6.4 Glottis Segmentation
We evaluate our segmentation model on the public glottis dataset BAGLS. We use Intersection over Union (IoU) as the metric, and the results are shown in Table. 5. From the table, we can see that the diffusion model can have a slightly better IoU performance than the traditional U-Net, By refining the U-Net results with the diffusion ones, the IoU performance can be further improved.
@@ 6.5 Vocal Fold Paralysis Detection
Table 6: The average classification performance (%) achieved using different audio-based methods, along with the standard deviation (reported in parentheses), is evaluated across experiments using 10-fold cross-validation on the SYSU-A dataset.
@@
Table 7: The average classification performance (%) achieved using different enhanced modules and modalities, along with the standard deviation (reported in parentheses), is evaluated across experiments using 10-fold cross-validation on the SYSU-A dataset.
@@
Audio+AGA - - 84.88($\pm$7.66) 75.57($\pm$9.57) 71.73($\pm$9.49) 87.41($\pm$8.14) 78.42($\pm$7.64) 87.41($\pm$8.14) 63.73($\pm$16.92) 75.57($\pm$9.57)
Audio+AGA - 85.82($\pm$4.71) 76.20($\pm$8.86) 73.67($\pm$9.92) 85.03($\pm$6.33) 78.49($\pm$6.57) 85.03($\pm$6.33) 67.37($\pm$17.47) 76.20($\pm$8.86)
Audio+AGA 87.04($\pm$3.69) 78.12($\pm$5.06) 73.32($\pm$6.03) 90.06($\pm$6.52) 80.52($\pm$4.00) 90.06($\pm$6.52) 66.18($\pm$11.35) 78.12($\pm$5.06)
6.6 Evaluation Metrics
To comprehensively evaluate the performance of our proposed system, we employed several widely recognized metrics, including accuracy, precision, sensitivity (recall), specificity, F-score, the mean of sensitivity and specificity, and the area under the receiver operating characteristic curve (ROC-AUC). Together, these metrics, as detailed in Table 7, highlight the robustness and clinical applicability of the proposed system.

### v4

@@ 1 Introduction
Vocal fold Paralysis (VFP) is a condition where one of the vocal folds fails to move properly, leading to voice changes, difficulty swallowing, and potential breathing problems [45, 14]. VFP can result from nerve damage due to surgery, injury, infection, or tumors, significantly impacting a patient's quality of life [36]. Accurate diagnosis of VFP is crucial, as it informs the appropriate medical or surgical intervention, which can restore vocal function, improve airway protection, and enhance overall patient outcomes. Clinicians often use laryngeal videostroboscopy to check the vocal fold vibration in details. Laryngeal videostroboscopy is a specialized diagnostic tool used to evaluate the function of the vocal folds and the larynx. The stroboscopic component is important because it allows for visualization of vocal fold vibrations, which are often too rapid to be observed directly by doctors [3]. Stroboscopy uses a light source that flashes at a slightly different frequency than the vocal fold vibration. This creates a slow-motion effect, enabling the clinician to see the individual phases of vocal fold vibration in great details.
With the advent of the artificial intelligence, deep learning methods are developed to extract useful parameters [39] to assist clinicians track the motion of vocal folds [9, 32, 13, 12, 49], and even make predictions [30, 7] or classifications [37] to help clinicians perform diagnosis. However, the diagnosis of vocal folds requires to inspect the complete phonation cycles of patients. Previous works default to using video frames or images that already contain the vocal folds, but in most of the time, the raw recordings from the endoscopic inspection often carry useless information other than patient's phonation cycles. For instance, no phonation cycles occur in the beginning of the inspection, as the laryngoscope is still finding the vocal folds. As a result, manual labeling and selection of video frames from raw endoscopic videos are required, making the diagnosis process time-consuming and inconvenient for clinicians.
@@
In this article, we introduce the Multimodal Laryngoscopic Video Analyzing System (MLVAS), a novel framework that leverages both audio and visual modalities to automatically process and analyze raw laryngeal videostroboscopic recordings. The proposed system comprises a multimodal front-end module, a feature extraction module, and a multimodal back-end classification module for detecting VFP. The front-end module identifies and extracts video segments that captures complete phonation cycles with clear vocal folds' movement from raw endoscopic laryngeal videos. Subsequently, MLVAS extracts both audio and video features from these segments. By leveraging these complementary modalities, the system employs a classifier in the back-end for VFP detection. The primary contributions of this work are as follows:
@@
    The system introduces Left and Right Vocal Fold Dynamics (LVFDyn and RVFDyn) to reflect the vocal fold's activity, enabling precise diagnosis of left- or right-sided VFP. Experimental findings highlight its effectiveness in diagnosing Unilateral Vocal Fold Paralysis (UVFP).
@@ 2.1 System Design
2.2 Audio Processing Module
To get a better examination of the status of the vocal folds, patients are often instructed to pronounce ``/i:/''. However, due to the physical constraints of the tongue during endoscopic procedures, the actual sound produced often resembles ``/:/'', which corresponds to the International Phonetic Alphabet (IPA) symbol. This sound pattern enables clinicians to detect the vibration of the vocal folds more effectively. Hence, by capturing the audio segments that pronounce ``/:/'', phonation cycles can be detected and analyzed. To accomplish this, we developed a KWS model that is used to detect audio keywords in a sentence, similar to those used in systems like ``Hey Siri'', ``Ok Google''. In our case, the KWS model is designed to identify instances of the keyword ``/:/''.
The overall pipeline is illustrated in Figure 2. Initially, the input audio is transformed into a STFT spectrogram, converting the time-domain signal into a time-frequency representation for more effective analysis. The spectrogram is then segmented into chunks along the time axis, with each chunk containing a fixed number of frames. During the training phase, these chunks are randomly selected from each audio clip to ensure a diverse set of training samples. This method enhances the model's generalization capabilities by incorporating spectrograms from various audio clips into a robust batch of training data.
@@ 2.2 Audio Processing Module
During inference, the input spectrogram is sliced into chunks using a sliding window mechanism. Each chunk is processed by the trained KWS model to generate decision results, as depicted in Figure 2. This KWS model can reliably detects the ``/:/'' vocalization, providing critical prior knowledge for the subsequent analysis.
@@ 2.3 Video Processing Module
Although the audio KWS model is employed to detect the vocalization video segment of the phonation ``/:/'', these detected video segments may not always successfully capture the vocal folds and glottis on the visual side. To address this issue, we train a vocal fold detection model, utilizing the famous object detection model structure YOLO-v5 [27] on the BAGLS dataset [11], to further refine the time masks.
As depicted in Figure 1, our pipeline integrates a strobing video extraction module alongside the vocal fold image object detection module. This module is utilized to select the strobing segments of laryngeal videos by analyzing hue, saturation, and value (HSV) of video frames. The HSV parameters represent type, intensity, and brightness of colors, respectively. Transitions marked by empty frames typically occur at the beginning and end of strobing segments. We employ a unit step function to identify and mark all empty frames with a zero value in color (illustrated by the yellow line in Figure 3). This technique segments the video into several continuous non-empty frame sequences. Within these non-empty segments, we examine the variations in HSV values. The variations is calculated as the total number of changes within a time range. Due to the rapid color-changing characteristic of strobing videos (as shown in Figure 3), we are able to identify the strobing segments by calculating the frequency of HSV fluctuations within each non-empty segment, selecting the segment with the highest number of fluctuations. Suppose $v_{t-1},v_{t},v_{t+1}$ stands for the value at time step t1,tt-1,t and $t+1$ respectively. Then, the frequency of HSV fluctuation within a time range of $T=[t_{0},t_{n}]$ is calculated as follows.
$F_{T}=\sum_{t=1}^{n-1}\frac{(v_{t}-v_{t-1})\cdot(v_{t+1}-v_{t})}{|(v_{t}-v_{t-1})\cdot(v_{t+1}-v_{t})|}$
@@ 4.2 Diffusion Model-based Refinement
Figure 5: The workflow of computing the angle deviation of left and right vocal folds. We use image#116 from BAGLS dataset as an example. The black background represents the unsegmented region, while the white area corresponds to the segmented glottis mask. Step (a) get the center point $C$ and bottom point $D$. Step (b), (c) and (d) fit the outline of vocal folds with a quadratic function, refining the bottom point to $D^{\prime}$. As a result, step (e) shows the refined midline $CD^{\prime}$. Finally, the angle deviation for both vocal folds can be computed, which are $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\angle L_{1}D^{\prime}C,\angle L_{2}D^{\prime}C,\angle R_{1}D^{\prime}C$ and $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\angle R_{2}D^{\prime}C$.
Table 4: Angle Deviation Extraction Algorithm and Corresponding Figures.
Algorithm Step Figure
1. Compute the center $C(x_{c},y_{c})$ from $U(x_{1},y_{1})$, $D(x_{2},y_{2})$, $L(x_{3},y_{3})$, $R(x_{4},y_{4})$. Figure 5(b)
2. Compute the function $f(x)$ passing through $C$ and $D$. Figure 5(b)
3. Find $N$ equidistant points $C_{1},C_{2},\dots,C_{N-1}$ of $f(x)$ on the line segment intercepted by the glottal mask. Figure 5(c)
4. For each $k=1,2,\dots,N-1$: 4.1 Compute the function $f_{k}^{\prime}(x)$ orthogonal to $f(x)$ passing through $C_{k}$. 4.2 Compute intersection points $L_{k}$ and $R_{k}$ between $f_{k}^{\prime}(x)$ and the glottis mask. Figure 5(c)
5. Rotate the coordinate system by $\gamma$ degrees to align $f(x)$ vertically. Figure 5(d)
6. Approximate a quadratic curve $q_{\gamma}(x)$ using the rotated intersection points $\{L_{\gamma,k}\}$ and $\{R_{\gamma,k}\}$. Figure 5(d)
7. Find the vertex point $D_{q}$ of $q_{\gamma}(x)$ and map it back to the original coordinate system. Figure 5(e)
8. Connect $C$ and $D_{q}$ to form the corrected midline $f^{*}(x)$, which intersects the glottis mask at $D^{\prime}$. Figure 5(e)
9. Connect $\{L_{k}\}$ and $\{R_{k}\}$ with $D^{\prime}$ to compute the glottal angles $\angle L_{k}D^{\prime}C$ and $\angle CD^{\prime}R_{k}$. Figure 5(f)
@@
4.3 Vocal Fold Dynamics (VFDyn) Extraction with Quadratic Fitting (QF)
To diagnose the laryngeal paralysis in a more detailed way, instead of AGA or GAW that measures the both left and right vocal folds, we evaluate the angle deviation for each vocal fold. As a result, we can provide a more comprehensive diagnosis of UVFP. The procedure for extracting glottal angles from a segmented glottis involves a systematic series of steps, detailed in Table 4, with visual representations of the intermediate results provided in Figure 5.
Figure 6: The pipeline of the multimodal VFP classification model.
Initially, the algorithm acquires the coordinates of the top, bottom, left, and right vertices from the glottis mask, computing their center point denoted as $U(x_{1},y_{1})$, $D(x_{2},y_{2})$, $L(x_{3},y_{3})$, $R(x_{4},y_{4})$, and $C(x_{c},y_{c})$ respectively (see Figure 5(b)). Subsequently, a line connecting points $C$ and $D$ is established, and a linear function $f(x)$ passing through $C$ and $D$ is computed.
Equidistant points $C_{1},C_{2},...,C_{N-1}$ are then positioned along the line segment intercepted by the glottal mask. For each $C_{k}$, orthogonal functions $f^{\prime}_{k}(x)$ to $f(x)$ are calculated, with their intersection points $L_{k}$ and $R_{k}$ with the glottis mask determined for $k\in[1,N-1]$ (see Figure 5(c)).
To ensure uniformity, the coordinate system is rotated by an angle $\gamma$, aligning all intersection points along the y-axis after rotation, denoted as $L_{\gamma,k}$ and $R_{\gamma,k}$. Leveraging these points, a quadratic curve $q_{\gamma}(x)$ is approximated in the rotated coordinate system (see Figure 5(d). The lowest point $D_{q}$ of $q_{\gamma}(x)$ is identified and mapped back to the original coordinate system (see Figure 5(e)).
Subsequently, a calibrated middle line $f^{*}(x)$ connecting points $C$ and $D_{q}$ is established, intersecting the glottis mask at point $D^{\prime}$. Finally, glottal angles are derived by connecting points $\{L_{k}\}_{k\in[1,N-1]}$ and $\{R_{k}\}_{k\in[1,N-1]}$ with $D^{\prime}$, yielding $\{\angle L_{k}D^{\prime}C\}_{k\in[1,N-1]}$ and $\{\angle CD^{\prime}R_{k}\}_{k\in[1,N-1]}$ respectively (see Figure 5(f)).
@@ 6.3 Keyword Spotting Model
Since the KWS model outputs a posterior score rather than a direct decision value, a threshold is necessary to determine the final result. Consequently, the threshold is selected based on the precision-recall curve on the training portion of the SYSU-B dataset, aiming for the highest F1-score, a balanced metric that considers both precision and recall, providing a comprehensive measure of classification performance. As a result, we pick a threshold of 0.38 as the classification boundary. If the posterior score output by the KWS model is higher than this value, the input chunk is regarded as ``/:/''. Otherwise, it is treated as other voices. Figure 7(a) and Figure 7(b) display the confusion matrix computed on the SYSU-B-test and SYSU-A dataset respectively. The accuracy, precision, recall and F-score are listed in Table 5. As is shown in Table 5, the performance of our audio KWS model are all above 90% on both the SYSU-A and SYSU-B.
@@ 6.4 Glottis Segmentation
We evaluate our segmentation model on the public glottis dataset BAGLS. We use Intersection over Union (IoU) as the metric, and the results are shown in Table 6. From the table, we can see that the diffusion model can have a slightly better IoU performance than the traditional U-Net, By refining the U-Net results with the diffusion ones, the IoU performance can be further improved.
@@ 6.5 Vocal Fold Paralysis Detection
6.6 Evaluation Metrics
To comprehensively evaluate the performance of our proposed system, we employed several widely recognized metrics, including accuracy, precision, sensitivity (recall), specificity, F-score, the mean of sensitivity and specificity, and the area under the receiver operating characteristic curve (ROC-AUC). Together, these metrics highlight the robustness and clinical applicability of the proposed system.
Table 7: The average classification performance (%) achieved using different audio-based methods, along with the standard deviation (reported in parentheses). The results are evaluated using 10-fold cross-validation on the SYSU-A dataset. All models operate on pre-extracted key video segments, which ensures the visibility of the larynx and includes at least one complete phonatory cycle.
@@ 6.6 Evaluation Metrics
Table 8: The average classification performance (%) achieved using different enhanced modules and modalities, along with the standard deviation (reported in parentheses), is evaluated across experiments using 10-fold cross-validation on the SYSU-A dataset. Audio: Dasheng-encoded audio features; VFDyn: Vocal fold dynamics extracted from key video segments.
@@
Audio+VFDyn - - 84.88($\pm$7.66) 75.57($\pm$9.57) 71.73($\pm$9.49) 87.41($\pm$8.14) 78.42($\pm$7.64) 87.41($\pm$8.14) 63.73($\pm$16.92) 75.57($\pm$9.57)
Audio+VFDyn - 85.82($\pm$4.71) 76.20($\pm$8.86) 73.67($\pm$9.92) 85.03($\pm$6.33) 78.49($\pm$6.57) 85.03($\pm$6.33) 67.37($\pm$17.47) 76.20($\pm$8.86)
Audio+VFDyn 87.04($\pm$3.69) 78.12($\pm$5.06) 73.32($\pm$6.03) 90.06($\pm$6.52) 80.52($\pm$4.00) 90.06($\pm$6.52) 66.18($\pm$11.35) 78.12($\pm$5.06)